# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ VIII-E Evaluation Metrics
As seen in Table IX, LLMs research predominantly utilizes traditional quantitative metrics such as bug reproduction rate, code coverage, precision, and recall, these metrics directly reflect the effectiveness and quality of test generation. In contrast, LLM-Based agents research not only focuses on quantitative metrics but also introduces qualitative evaluations, such as improvements through conversational interactions and the collaborative effects of multi-agent systems. This diversified evaluation approach provides a more comprehensive reflection of the system’s practical application effects. From the task perspectives, LLMs are more inclined to single task processing, such as generating test sets and considering the coverage of generated test sets. However, because of the expansion of agents framework, LLM-based agents often tend to use the generated test sets to evaluate whether vulnerabilities can be found to achieve a more ideal practicality. From a design perspective, LLM systems are relying on prompt engineering and the generative capabilities of the models themselves, their evaluation metrics are also mainly focused on the quality and effectiveness of the model outputs, also their evaluation metrics include the collaborative effects and efficiency within the system, such as improving Pass@1 and Complete@1 rates through multi-agent collaboration. Overall, LLMs are more suited for rapid test generation and evaluation for specific tasks, with evaluation metrics directly reflecting the generation’s effectiveness and quality. LLM-Based Agents excel in handling complex and diversified tasks, achieving higher system efficiency and effectiveness through multi-agent collaboration and iterative optimization.
TABLE IX: Evaluation Metrics in Software Test Generation
@@ IX Software Security and Maintenance
In the software engineering, software security and maintenance is a popular area for the application of LLMs, primarily aimed at enhancing the security and stability of software systems through existing technologies to meet the needs of users and developers. These models provide promise methods of vulnerability detection and repair, while also enabling automated security testing and innovative maintenance processes. The application of LLMs in software security and maintenance encompasses several aspects, including vulnerability detection, automatic repair, penetration testing, and system robustness evaluation. Compared to traditional methods, LLMs leverage natural language processing and generation technologies to understand and generate complex code and security policies, thereby automating detection and repair tasks. For example, LLMs can accurately identify potential vulnerabilities by analyzing code structures and contextual information and generate corresponding repair suggestions which improves the efficiency and accuracy of vulnerability recovery.
@@ IX-A LLMs Tasks
In the field of software security and maintenance, research on LLMs can be categorized into three main areas: vulnerability detection, automatic repair, and penetration testing, along with some evaluation studies. The collected papers reviewed on LLMs in these domain illustrate their diverse applications and potential.
IX-A1 Program Vulnerability
In the domain of vulnerability detection, researchers have fine-tuned LLMs to enhance the accuracy of source code vulnerability detection. [146] aims to investigate the potential of applying LLMs to the task of vulnerability detection in source code and to determine if the performance limits of CodeBERT-like models are due to their limited capacity and code understanding ability. The study fine-tuned the WizardCoder model (an improved version of StarCoder) and compared its performance with the ContraBERT model on balanced and unbalanced datasets. The experimental results showed that WizardCoder outperformed ContraBERT in both ROC AUC and F1 scores, significantly improving Java function vulnerability detection performance, which achieved the state-of-art performance at that time by improving ROC AUC from 0.66 in CodeBERT to 0.69.
There are study mainly explored the applications of pure LLMs without any framework architecture in vulnerability detection, uncovering current challenges. [147] evaluated only the performance of ChatGPT and GPT-3 models in detecting vulnerabilities in Java code, the study compared text-davinci-003 (GPT-3) and gpt-3.5-turbo against a baseline virtual classifier in binary and multi-label classification tasks. The experimental results showed that while text-davinci-003 and gpt-3.5-turbo had high accuracy and recall rates in binary classification tasks, their AUC (Area Under Curve) scores were only 0.51, indicating performance equivalent to random guessing. In multi-label classification tasks, GPT-3.5-turbo and text-davinci-003 did not significantly outperform the baseline virtual classifier in overall accuracy and F1 scores. These findings indicate that the earlier model like GPT-3 has limited capabilities in practical vulnerability detection tasks, suggesting the need for further research and model optimization to improve their performance in real-world applications, fine-tuning and optimizing LLMs can significantly enhance their performance in source code vulnerability detection, However, these models still face many challenges in practical applications, requiring further research and technological improvements to enhance their real-world effectiveness and reliability.
@@ IX-A1 Program Vulnerability
IX-A2 Automating Program Repair
@@
In the domain of software security and maintenance, LLMs have not only been applied to vulnerability detection but also extensively used for automating program repair. One study proposed using Round-Trip Translation (RTT) for automated program repair, researchers translated defective code into another language and then back to the original language to generate potential patches. The study used various language models and benchmarks to evaluate RTT’s performance in APR. The experiments explored how RTT performs when using programming languages as an intermediate representation, how RTT performs when using natural language (English) as an intermediate representation, and what qualitative trends can be observed in the patches generated by RTT. Three measurement standards and eight models were used in the experiments, the results showed that the RTT method achieved significant repair effects on multiple benchmarks, particularly excelling in terms of compilation and feasibility [151]. Similarly, in automated program repair, [145] introduced several innovative methods. For example, NAVRepair specifically targets C/C++ code vulnerabilities by combining node type information and error types. Due to the unique pointer operations and memory management issues in C/C++, this language poses complexities. The framework uses Abstract Syntax Trees (ASTs) to extract node type information and combines it with CWE-derived vulnerability templates to generate targeted repair suggestions, the study evaluated NAVRepair on several popular LLMs (ChatGPT, DeepSeek Coder, and Magicoder) to demonstrate its effectiveness in improving code vulnerability repair performance. The results showed that NAVRepair achieved state-of-art performance in C/C++ program repair task, which improved repair accuracy by 26% compared to existing methods.
@@
In order to address the two main limitations of existing fine-tuning methods for LLM-based program repair, which is the lack of reasoning about the logic behind code changes and high computational costs associated with fine-tuning large datasets. [152] introduced the MOREPAIR framework, this framework improve the performance of LLMs in automated program repair (APR) by simultaneously optimizing syntactic code transformations and the logical reasoning behind code changes, the study used techniques to enhance fine-tuning efficiency, such as QLoRA (Quantized Low-Rank Adaptation) [153] to reduce memory requirements and NEFTune (Noisy Embedding Fine-Tuning) [154] to prevent overfitting during the fine-tuning process. The experiments evaluated MOREPAIR on four open-source LLMs of different sizes and architectures (CodeLlama-13B, CodeLlama-7B, StarChat-alpha, and Mistral-7B) using two benchmarks, evalrepair-C++ and EvalRepair-Java. The results indicated that CodeLlama improved by 11% and 8% on the first 10 repair suggestions for evalrepair-C++ and EvalRepair-Java respectively. Another study introduced the PyDex system, which automatically repairs syntax and semantic errors in introductory Python programming assignments using LLMs, the system combines multimodal prompts and iterative querying methods to generate repair candidates and uses few-shot learning to improve repair accuracy. PyDex was evaluated on 286 real student programs from an introductory Python programming course and compared against three baselines. The results showed that PyDex significantly improved the repair rate and effectiveness compared to existing baselines [155].
@@
IX-A3 Penetration Testing
LLMs can also be applied in the field of penetration testing, where they are used to enhance the efficiency and effectiveness of automated penetration testing. Although not as frequently studied as vulnerability detection and automated repair, this review includes two relevant papers. [143] investigates the development and evaluation of an LLM-driven automatic penetration testing tool PENTESTGPT. The main purpose of this study is to evaluate the performance of LLMs in practical penetration testing tasks and address the issue of context loss during the penetration testing process, the paper introduces three self-interaction modules of PENTESTGPT (reasoning, generation, and parsing) and provides empirical research based on benchmarks involving 13 targets and 182 sub-tasks. It compares the penetration testing performance of GPT-3.5, GPT-4, and Bard. The experimental results show that PENTESTGPT’s task completion rate is 228.6% higher than GPT-3.5 and 58.6% higher than GPT-4, this study demonstrates the potential of LLMs in automated penetration testing, helping to identify and resolve security vulnerabilities, thereby enhancing the security and robustness of software systems.
@@ IX-B LLM-based Agents Task
LLM-based Agents primarily appied in areas such as autonomous decision-making, task-specific optimization, and multi-agent collaboration, these frameworks showcasing their strong potential in proactive defense.  [159] aims to address the limitations of existing debugging methods that treat the generated program as an indivisible entity. By segmenting the program into basic blocks and verifying the correctness of each block based on task descriptions, the proposed method LDB (Large Language Model Debugger) provide a more detailed and effective debugging tool that closely reflects human debugging practices. The study’s experiments covered testing LDB on several benchmarks and compared with baseline models without a debugger and those using traditional debugging methods (self-debugging with explanations and traces). LDB’s accuracy increased from a baseline of 73.8% to 82.9% on the HumanEval benchmark, an improvement of 9.1%. In the domain of vulnerability detection, researchers have enhanced detection accuracy by combining Role-Based Access Control (RBAC) practices with deep learning of complex code structures.
@@
[160] addresses the challenge of automatically and appropriately repairing access control (AC) vulnerabilities in smart contracts. The innovation of this paper lies in combining mined RBAC practices with LLMs to create a context-aware repair framework for AC vulnerabilities. The model primarily uses GPT-4, enhanced by a new method called ACFIX, which mines common RBAC practices from existing smart contracts and employs a Multi-Agent Debate (MAD) mechanism to verify the generated patches through debates between generator and verifier agents to ensure correctness. Experimental results show that ACFIX successfully repaired 94.92% of access control vulnerabilities, significantly outperforming the baseline GPT-4’s 52.54%. Another application in smart contracts  [161], this paper introduces a two-stage adversarial framework, GPTLENS, which improves vulnerability detection accuracy through generation and discrimination phases. GPTLENS achieved a 76.9% success rate in detecting smart contract vulnerabilities, better than the 38.5% success rate of traditional methods. Another study, [109] investigates the use of GPT-4 to automatically exploit disclosed but unpatched vulnerabilities, the experiments showed that the LLM-based agent achieved an 87% success rate in exploiting vulnerabilities when provided with CVE descriptions. Finally another LLM-based agent application in the penetration test,  [107] employs GPT-3.5 to assist penetration testers by automating high-level task planning and low-level vulnerability discovery, thereby enhancing penetration testing capabilities. The experiments demonstrated successful automation of multiple stages of penetration testing, including high-level strategy formulation and low-level vulnerability discovery, showcasing the effectiveness of LLMs in penetration testing.
In the field of software repair by multi-agents collaborations, [162] proposes a dual-agent framework that enhances the automation and accuracy of repairing declarative specifications through iterative prompt optimization and multi-agent collaboration. The researcher compare the effectiveness of the LLM-based repair pipeline with several state-of-the-art Alloy APR techniques (ARepair, ICEBAR, BeAFix, and ATR). In the result, framework repaired 231 defects in the Alloy4Fun benchmark which surpassing the 278 defects repaired by traditional tools. In  [142], developed and evaluated an automated debugging framework named FixAgent, which improves fault localization, repair generation, and error analysis through an LLM-based multi-agent system. Although this research primarily focuses on automated debugging, incorporating elements like fault localization and automated program repair (APR), it intersects with test generation, particularly in the validation phase for testing bug fixes. The study evaluates FixAgent’s performance on the Codeflaws, QuixBugs, and ConDefects datasets, comparing it to 16 baseline methods, including state-of-the-art APR tools and LLMs. Experimental results show that FixAgent fixed 78 out of 79 bugs in the QuixBugs dataset, including 9 never-before-fixed bugs. In the Codeflaws dataset, FixAgent fixed 3982 out of 2780 defects, with a correctness rate of 96.5%. The framework includes specialized agents responsible for localization, repair, and analysis tasks and uses the rubber duck debugging principle. FixAgent demonstrates the powerful capabilities of LLMs in automated debugging, improving the performance of existing APR tools and LLMs which can be considered as state-of-art framework in APR by LLM-based agent.
[46] introduces an automated program repair agent named RepairAgent, this agent can dynamically generates prompts and integrates tools to automatically fix software bugs. This researcher also address the limitations of current LLM-based repair techniques, which typically involve fixed prompts or feedback loops that do not allow the model to gather comprehensive information about the bug or code. RepairAgent is a LLM-based agent designed to alternately collect information about the bug, gather repair ingredients, and validate the repairs, similar to how human developers fix bugs. RepairAgent achieved impressive result by overall repaired 186 bugs in the Defects4J benchmark, with 164 being correctly repaired outperforming existing repair techniques achieved the state-of-art performances.
@@
In the realm of software security, researchers have combined LLM and security engineering models to improve security analysis and design processes. [163] aims to propose a complex hybrid strategy to ensure the reliability and security of software systems, this involves a concept-guided approach where LLM-based agents interact with system model diagrams to perform tasks related to safety analysis. [108] introduces the TrustLLM framework which increase the accuracy and interpretability of smart contract auditing by customizing LLM capabilities to the specific requirements of smart contract code. This paper conducts experiments on a balanced dataset comprising 1,734 positive samples and 1,810 negative samples, comparing TrustLLM with other models such as CodeBERT, GraphCodeBERT, and several versions of GPT and CodeLlama. TrustLLM achieves an F1 score of 91.21% and an accuracy of 91.11% which outperforming other models. Beyond software-level security design, LLMs can also be integrated into autonomous driving systems. [164] which has already been discussed in the IV.

### v2

@@ VIII-E Evaluation Metrics
As seen in Table X, LLMs research predominantly utilizes traditional quantitative metrics such as bug reproduction rate, code coverage, precision, and recall, these metrics directly reflect the effectiveness and quality of test generation. In contrast, LLM-Based agents research not only focuses on quantitative metrics but also introduces qualitative evaluations, such as improvements through conversational interactions and the collaborative effects of multi-agent systems. This diversified evaluation approach provides a more comprehensive reflection of the system’s practical application effects. From the task perspectives, LLMs are more inclined to single task processing, such as generating test sets and considering the coverage of generated test sets. However, because of the expansion of agents framework, LLM-based agents often tend to use the generated test sets to evaluate whether vulnerabilities can be found to achieve a more ideal practicality. From a design perspective, LLM systems are relying on prompt engineering and the generative capabilities of the models themselves, their evaluation metrics are also mainly focused on the quality and effectiveness of the model outputs, also their evaluation metrics include the collaborative effects and efficiency within the system, such as improving Pass@1 and Complete@1 rates through multi-agent collaboration. Overall, LLMs are more suited for rapid test generation and evaluation for specific tasks, with evaluation metrics directly reflecting the generation’s effectiveness and quality. LLM-Based Agents excel in handling complex and diversified tasks, achieving higher system efficiency and effectiveness through multi-agent collaboration and iterative optimization.
TABLE X: Evaluation Metrics in Software Test Generation
@@
DynaMOSA benchmark
Defects4J.
Correctness, Readability,
Code Coverage, Bug Detection
Vargha-Delaney A12.
No
[169]
@@ IX Software Security and Maintenance
In software engineering, software security and maintenance is a popular area for the application of LLMs, primarily aimed at enhancing the security and stability of software systems through existing technologies to meet the needs of users and developers. These models provide promising methods of vulnerability detection and repair, while also enabling automated security testing and innovative maintenance processes. The application of LLMs in software security and maintenance encompasses several aspects, including vulnerability detection, automatic repair, penetration testing, and system robustness evaluation. Compared to traditional methods, LLMs leverage natural language processing and generation technologies to understand and generate complex code and security policies, thereby automating detection and repair tasks. For example, LLMs can accurately identify potential vulnerabilities by analyzing code structures and contextual information and generate corresponding repair suggestions which improves the efficiency and accuracy of vulnerability recovery.
@@ IX-A LLMs Tasks
In the field of software security and maintenance, research on LLMs can be categorized into three main areas: vulnerability detection, automatic repair, and penetration testing, along with some evaluation studies. The collected papers reviewed on LLMs in these domains illustrate their diverse applications and potential.
Static Vulnerability Detection From Prompting to Structural Learning. In the domain of vulnerability detection, researchers have fine-tuned LLMs to enhance the accuracy of source code vulnerability detection. [174] aims to investigate the potential of applying LLMs to the task of vulnerability detection in source code and to determine if the performance limits of CodeBERT-like models are due to their limited capacity and code understanding ability. The study fine-tuned the WizardCoder model (an improved version of StarCoder) and compared its performance with the ContraBERT model on balanced and unbalanced datasets. The experimental results showed that WizardCoder outperformed ContraBERT in both ROC AUC and F1 scores, significantly improving Java function vulnerability detection performance, which achieved the state-of-art performance at that time by improving ROC AUC from 0.66 in CodeBERT to 0.69.
There are study mainly explored the applications of pure LLMs without any framework architecture in vulnerability detection, uncovering current challenges. [175] evaluated only the performance of ChatGPT and GPT-3 models in detecting vulnerabilities in Java code, the study compared text-davinci-003 (GPT-3) and gpt-3.5-turbo against a baseline virtual classifier in binary and multi-label classification tasks. The experimental results showed that while text-davinci-003 and gpt-3.5-turbo had high accuracy and recall rates in binary classification tasks, their AUC (Area Under Curve) scores were only 0.51, indicating performance equivalent to random guessing. In multi-label classification tasks, GPT-3.5-turbo and text-davinci-003 did not significantly outperform the baseline virtual classifier in overall accuracy and F1 scores. These findings indicate that the earlier model like GPT-3 has limited capabilities in practical vulnerability detection tasks, suggesting the need for further research and model optimization to improve their performance in real-world applications, fine-tuning and optimizing LLMs can significantly enhance their performance in source code vulnerability detection, However, these models still face many challenges in practical applications, requiring further research and technological improvements to enhance their real-world effectiveness and reliability. Recent work also emphasizes the importance of fair evaluation settings and advanced prompt engineering strategies in LLM-based program repair.  [176] constructs a new benchmark dataset, EvalGPTFix, consisting of 151 real-world Java bug-fix pairs collected from 2023 AtCoder [^8] https://atcoder.jp/contests/ahc030 contests to eliminate data leakage risks inherent in previous benchmarks such as Defects4J. Through empirical studies, the paper demonstrates that ChatGPT, using only a basic prompt, can correctly repair 109 bugs (72.19% recall), significantly outperforming state-of-the-art LLMs such as CodeT5 [177] and PLBART by 27.5% and 62.4% respectively, This work highlights the critical role of prompt design and interaction in maximizing LLM capabilities for program repair.
@@
Empirical Evaluation of Web Vulnerabilities in LLM-Generated Code. While many vulnerability detection efforts focus on static code analysis or fine-tuning models, an emerging direction is the large-scale empirical evaluation of LLM-generated code for real-world deployment readiness and security risks. One recent study presents a comprehensive evaluation of 2,500 GPT-4-generated PHP web applications, analyzing their vulnerability using hybrid techniques including dynamic scanning, static analysis, and manual code audit [181]. The authors constructed the ChatPHP-DB dataset, which contains PHP code and corresponding SQL scripts automatically generated by LLMs. Vulnerability scanning was conducted on web services deployed in Docker environments, focusing on common web attack surfaces such as SQL Injection, Stored XSS, Reflected XSS, and Insecure File Upload. The authors highlighted that many vulnerabilities stemmed from overly simplistic coding patterns and missing sanitization, and that LLMs often fail to internalize secure coding practices, especially in zero-shot settings.
@@
Generalized and Specialized Program Repair. In the domain of software security and maintenance, LLMs have not only been applied to vulnerability detection but also extensively used for automating program repair. One study proposed using Round-Trip Translation (RTT) for automated program repair, researchers translated defective code into another language and then back to the original language to generate potential patches. The study used various language models and benchmarks to evaluate RTT’s performance in APR. The experiments explored how RTT performs when using programming languages as an intermediate representation, how RTT performs when using natural language (English) as an intermediate representation, and what qualitative trends can be observed in the patches generated by RTT. Three measurement standards and eight models were used in the experiments, the results showed that the RTT method achieved significant repair effects on multiple benchmarks, particularly excelling in terms of compilation and feasibility [182]. Similarly, in automated program repair, [173] introduced several innovative methods. For example, NAVRepair specifically targets C/C++ code vulnerabilities by combining node type information and error types. Due to the unique pointer operations and memory management issues in C/C++, this language poses complexities. The framework uses Abstract Syntax Trees (ASTs) to extract node type information and combines it with CWE-derived vulnerability templates to generate targeted repair suggestions, the study evaluated NAVRepair on several popular LLMs (ChatGPT, DeepSeek Coder, and Magicoder) to demonstrate its effectiveness in improving code vulnerability repair performance. The results showed that NAVRepair achieved state-of-art performance in C/C++ program repair task, which improved repair accuracy by 26% compared to existing methods.
@@
In order to address the two main limitations of existing fine-tuning methods for LLM-based program repair—namely, the lack of reasoning about the logic behind code changes and the high computational costs associated with fine-tuning large datasets— [183] introduced the MOREPAIR framework, this framework improve the performance of LLMs in automated program repair (APR) by simultaneously optimizing syntactic code transformations and the logical reasoning behind code changes, the study used techniques to enhance fine-tuning efficiency, such as QLoRA (Quantized Low-Rank Adaptation) [184] to reduce memory requirements and NEFTune (Noisy Embedding Fine-Tuning) [185] to prevent overfitting during the fine-tuning process. The experiments evaluated MOREPAIR on four open-source LLMs of different sizes and architectures (CodeLlama-13B, CodeLlama-7B, StarChat-alpha, and Mistral-7B) using two benchmarks, evalrepair-C++ and EvalRepair-Java. The results indicated that CodeLlama improved by 11% and 8% on the first 10 repair suggestions for evalrepair-C++ and EvalRepair-Java respectively. Another study introduced the PyDex system, which automatically repairs syntax and semantic errors in introductory Python programming assignments using LLMs, the system combines multimodal prompts and iterative querying methods to generate repair candidates and uses few-shot learning to improve repair accuracy. PyDex was evaluated on 286 real student programs from an introductory Python programming course and compared against three baselines. The results showed that PyDex significantly improved the repair rate and effectiveness compared to existing baselines [186].
@@
Penetration Testing. LLMs can also be applied in the field of penetration testing, where they are used to enhance the efficiency and effectiveness of automated penetration testing. Although not as frequently studied as vulnerability detection and automated repair, this review includes two relevant papers. [171] investigates the development and evaluation of an LLM-driven automatic penetration testing tool PENTESTGPT. The main purpose of this study is to evaluate the performance of LLMs in practical penetration testing tasks and address the issue of context loss during the penetration testing process, the paper introduces three self-interaction modules of PENTESTGPT (reasoning, generation, and parsing) and provides empirical research based on benchmarks involving 13 targets and 182 sub-tasks. It compares the penetration testing performance of GPT-3.5, GPT-4, and Bard. The experimental results show that PENTESTGPT’s task completion rate is 228.6% higher than GPT-3.5 and 58.6% higher than GPT-4, this study demonstrates the potential of LLMs in automated penetration testing, helping to identify and resolve security vulnerabilities, thereby enhancing the security and robustness of software systems.
@@ IX-B LLM-based Agents Task
Autonomous Debugging and Program Repair Agents. LLM-based Agents primarily appied in areas such as autonomous decision-making, task-specific optimization, and multi-agent collaboration, these frameworks showcasing their strong potential in proactive defense.  [190] aims to address the limitations of existing debugging methods that treat the generated program as an indivisible entity. By segmenting the program into basic blocks and verifying the correctness of each block based on task descriptions, the proposed method LDB (Large Language Model Debugger) provide a more detailed and effective debugging tool that closely reflects human debugging practices. The study’s experiments covered testing LDB on several benchmarks and compared with baseline models without a debugger and those using traditional debugging methods (self-debugging with explanations and traces). LDB’s accuracy increased from a baseline of 73.8% to 82.9% on the HumanEval benchmark, an improvement of 9.1%. In the domain of vulnerability detection, researchers have enhanced detection accuracy by combining Role-Based Access Control (RBAC) practices with deep learning of complex code structures. In  [170], developed and evaluated an automated debugging framework named FixAgent, which improves fault localization, repair generation, and error analysis through an LLM-based multi-agent system. Although this research primarily focuses on automated debugging, incorporating elements like fault localization and automated program repair (APR), it intersects with test generation, particularly in the validation phase for testing bug fixes. The study evaluates FixAgent’s performance on the Codeflaws, QuixBugs, and ConDefects datasets, comparing it to 16 baseline methods, including state-of-the-art APR tools and LLMs. Experimental results show that FixAgent fixed 78 out of 79 bugs in the QuixBugs dataset, including 9 never-before-fixed bugs. In the Codeflaws dataset, FixAgent fixed 3982 out of 2780 defects, with a correctness rate of 96.5%. The framework includes specialized agents responsible for localization, repair, and analysis tasks and uses the rubber duck debugging principle. FixAgent demonstrates the powerful capabilities of LLMs in automated debugging, improving the performance of existing APR tools and LLMs which can be considered as state-of-art framework in APR by LLM-based agent. In the field of software repair by multi-agents collaborations, [191] proposes a dual-agent framework that enhances the automation and accuracy of repairing declarative specifications through iterative prompt optimization and multi-agent collaboration. The researcher compare the effectiveness of the LLM-based repair pipeline with several state-of-the-art Alloy APR techniques (ARepair, ICEBAR, BeAFix, and ATR). In the result, framework repaired 231 defects in the Alloy4Fun benchmark which surpassing the 278 defects repaired by traditional tools. [51] introduces an automated program repair agent named RepairAgent, this agent can dynamically generates prompts and integrates tools to automatically fix software bugs. This researcher also address the limitations of current LLM-based repair techniques, which typically involve fixed prompts or feedback loops that do not allow the model to gather comprehensive information about the bug or code. RepairAgent is a LLM-based agent designed to alternately collect information about the bug, gather repair ingredients, and validate the repairs, similar to how human developers fix bugs. RepairAgent achieved impressive result by overall repaired 186 bugs in the Defects4J benchmark, with 164 being correctly repaired outperforming existing repair techniques achieved the state-of-art performances. Another notable work in autonomous program improvement is AutoCodeRover [192], which introduces a fully automated LLM-based agent framework for resolving GitHub issues through structured program understanding and modification. Different from previous agent frameworks that treat source code as flat files, AutoCodeRover leverages program representations such as Abstract Syntax Trees (AST) and introduces a two-stage architecture: a context retrieval agent and a patch generation agent. The context retrieval agent interacts with a set of designed APIs to iteratively collect relevant code context, while the patch generation agent synthesizes and validates patches based on the gathered information. This work marks a significant step toward scalable, low-cost, and semantically-aware autonomous software engineering. MAGIS [193] addresses repository-level issue resolution, which introduces a novel LLM-based multi-agent framework to tackle the complex task of resolving GitHub issues. MAGIS simulates a realistic team collaboration setting by deploying four specialized agents—Manager, Repository Custodian, Developer, and Quality Assurance (QA) Engineer. These agents cooperate through planning, code localization, iterative development, and automated review, closely mimicking human workflows like GitHub Flow.
@@
Access Control and Smart Contract Security. [194] addresses the challenge of automatically and appropriately repairing access control (AC) vulnerabilities in smart contracts. The innovation of this paper lies in combining mined RBAC practices with LLMs to create a context-aware repair framework for AC vulnerabilities. The model primarily uses GPT-4, enhanced by a new method called ACFIX, which mines common RBAC practices from existing smart contracts and employs a Multi-Agent Debate (MAD) mechanism to verify the generated patches through debates between generator and verifier agents to ensure correctness. Experimental results show that ACFIX successfully repaired 94.92% of access control vulnerabilities, significantly outperforming the baseline GPT-4’s 52.54%. Another application in smart contracts  [195], this paper introduces a two-stage adversarial framework, GPTLENS, which improves vulnerability detection accuracy through generation and discrimination phases. GPTLENS achieved a 76.9% success rate in detecting smart contract vulnerabilities, better than the 38.5% success rate of traditional methods. Another study, [137] introduces the TrustLLM framework which increase the accuracy and interpretability of smart contract auditing by customizing LLM capabilities to the specific requirements of smart contract code. This paper conducts experiments on a balanced dataset comprising 1,734 positive samples and 1,810 negative samples, comparing TrustLLM with other models such as CodeBERT, GraphCodeBERT, and several versions of GPT and CodeLlama. TrustLLM achieves an F1 score of 91.21% and an accuracy of 91.11% which outperforming other models.
LLMs for Offensive Security and Vulnerability Exploitation. [138] investigates the use of GPT-4 to automatically exploit disclosed but unpatched vulnerabilities, the experiments showed that the LLM-based agent achieved an 87% success rate in exploiting vulnerabilities when provided with CVE descriptions. Finally another LLM-based agent application in the penetration test,  [136] employs GPT-3.5 to assist penetration testers by automating high-level task planning and low-level vulnerability discovery, thereby enhancing penetration testing capabilities. The experiments demonstrated successful automation of multiple stages of penetration testing, including high-level strategy formulation and low-level vulnerability discovery, showcasing the effectiveness of LLMs in penetration testing.
@@
LLMs for Security Engineering and System-level Assurance. In the realm of software security, researchers have combined LLM and security engineering models to improve security analysis and design processes. [196] aims to propose a complex hybrid strategy to ensure the reliability and security of software systems, this involves a concept-guided approach where LLM-based agents interact with system model diagrams to perform tasks related to safety analysis. Beyond software-level security design, LLMs can also be integrated into autonomous driving systems. [80] which has already been discussed in the IV.